OpenStudy (anonymous):

by foil method: (7x+y) (2x-3y)

OpenStudy (anonymous):

first = (7x)(2x) = 14x^2 outter = (7x)(-3y) = -21xy inner = (y)(2x) = 2xy last = (y)(-3y) = -3y^2 then add all: 14x^2 - 20xy - 3y^2

jhonyy9 (jhonyy9):

there is -19xy and not -21xy ATTENTION please !!! because -21xy+2xy=-19xy

jhonyy9 (jhonyy9):

hope that is understandably

OpenStudy (anonymous):

\[F: 7x \times 2x = 14x^2\] \[O: 7x \times -3y = -21xy\] \[I: y \times 2x = 2xy\] \[L: y \times -3xy^2\] \[FOIL: 14x^2-19xy-3y^2\]